A person has 8 children of them he takes 3 at a time to a circus. He does not take


the same three children twice to the circus. How many times he will have to go


to circus to ensure that every three children have seen the circus together? In


this case find the number of times a particular child has visited the circus.





From 8 men and 4 women and team of 5 is to be formed. In how many ways can


this be done so as to include at least one woman?

1
Expert's answer
2021-12-28T11:34:46-0500

No 3 same children taken again required number = "^{8}C_3=\frac{8!}{3!.5!}=56"


Total number of men =8 

Total number of women =4 

Team = 5 persons → at least 1 women 

Case 1→1 women + 4 men ⇒"^4C_1 \times ^{8}C_4=4\times70=280"

Case 2→2 women + 3 men ⇒ "^4C_2 \times ^{8}C_3=6\times56=336"

Case 3→ 3 women + 2 men ⇒"^4C_3 \times ^{8}C_2=4\times28=112"

Case 4→ 4 women + 1 men ⇒"^4C_4 \times ^{8}C_1=1\times8=8"

Total ways "=280+336+112+8=736"